A, B & C form the vertices of a triangle.

 CAB = 90°, ∠

 ABC = 36° and AC = 8.8.

Calculate the length of AB rounded to 3 SF.



1
Expert's answer
2022-05-03T16:06:01-0400

Given triangle has right angle CAB.


So AB, AC and ∠ABC form the equation:


tan(∠ABC) = AC/AB.


From this equation we can find AB.


AB = AC / tan(∠ABC) = 8.8 / tan(36°) = 8.8 / 0.72654 = 12.112.


Answer: AB = 12.112.